Simplify:
sqrt45sqrt2000
I can solve up to:
sqrt45 => sqrt9*5 => 3sqrt5
When I get up to sqrt2000,
I'm having trouble in deciding what numbers to use to break it down.

Answer by stanbon(75887)

Simplify:
sqrt45sqrt2000
I can solve up to:
sqrt45 => sqrt9*5 => 3sqrt5
When I get up to sqrt2000,
I'm having trouble in deciding what numbers to use to break it down.
-----
2000 = 100*20 = 100*4*5
----
sqrt(2000) = sqrt(100*4*5) = 10*2*sqrt(5) = 20*sqrt(5)
------------
Cheers,
Stan H.
---------